OpenStudy (anonymous):

What is the volume occupied by 0.112mol of helium gas at a pressure of 0.96atm and a temperature of 312K

OpenStudy (anonymous):

Use the formula pV=nRT, where P is pressure in Pascals and take note that 1atm=101000Pa so 0.96=96960Pa. Temperature in Kelvin, n are moles and R is the contant gas value which is 8.31. Making n subject of the formula is V=nRT/P= 0.112*8.31*312/96960=0.002 m^3